OpenStudy (anonymous):

@sadellite73 please help me on this

OpenStudy (anonymous):

Compare the functions shown below: f(x) = 2 sin (3x + π) − 2 g(x) = (x − 3)2 − 1 h(x) x y -2 3 -1 -2 0 -5 1 -6 2 -5 3 -2 4 3 Which function has the smallest minimum y-value? f(x) g(x) h(x) Both g(x) and h(x) have the same minimum y-value.

OpenStudy (anonymous):

wrong one

OpenStudy (anonymous):

Compare the functions shown below: f(x) = 2 sin (3x + π) − 2 g(x) = (x − 3)2 − 1 h(x) x y -2 3 -1 -2 0 -5 1 -6 2 -5 3 -2 4 3 Which function has the smallest minimum y-value? f(x) g(x) h(x) Both g(x) and h(x) have the same minimum y-value.

OpenStudy (dtan5457):

@satellite73

OpenStudy (anonymous):

for h(x) the smallest value of y is -6, right?

OpenStudy (anonymous):

for g(x) the vertex of upward parabola is (3,-1), hence the smallest value of y is -1, right?

OpenStudy (anonymous):

for f(x) , the sin function has the amplitude |2|, hence the smallest value of y is -2. However, the whole thing -2 at the end of the right hand side forces you down the whole graph 2 units. So, the smallest value of y of f(x) is -4 Now, you have 3 of them, compare and make conclusion by yourself.

OpenStudy (anonymous):

i hate ur mom

OpenStudy (anonymous):

hahaha.... I think because my mom has a very cute baby and your mom does not.